OpenStudy (anonymous):

cot((x - (pi/2)) = -tan x

OpenStudy (anonymous):

\[ \cot (x-\pi/2) = \frac{\cos(x-\pi/2)}{\sin(x-\pi/2)} \]

OpenStudy (anonymous):

what do i do with this?

OpenStudy (anonymous):

Remember \[ \cos(\pi/2-x) = \sin(x) \]

OpenStudy (anonymous):

ok so is that the answer? or do i have to do something else?

OpenStudy (anonymous):

You have to keep manipulating it until it becomes what you want it to be.

OpenStudy (anonymous):

so do i foil it?

OpenStudy (anonymous):

Okay so \(x-\pi/2 = -(\pi/2-x)\)

OpenStudy (anonymous):

\[ \cos(-x) = \cos(x) \quad \sin(-x) = -\sin(x) \]